# Log-Scale Quantization in Distributed First-Order Methods: Gradient-based Learning from Distributed Data

### v2

@@ III-C Discussions
For convergence we only need $(\mathbf{1}_{n}\otimes I_{p})^{\top}H(\mathbf{1}_{n}\otimes I_{p})\succ 0$ (with $H:=\mbox{diag}[\nabla^{2}f_{i}(\mathbf{x}_{i})]$), i.e., strong-convexity of the global objective function $F(\mathbf{x})$. Thus, the local cost functions $f_{i}(\mathbf{x}_{i})$ might be non-convex, i.e., $\nabla^{2}f_{i}(\mathbf{x}_{i})$ is not necessarily positive while their summation is positive to ensure that $F(\mathbf{x})$ is strongly convex. An academic example of such a case is given in the simulation. Therefore, the convexity assumption in this paper is more relaxed as compared to many existing literature. Note that the log-scale quantization in this work is independent of Assumption 2.
@@
From Theorem 1, the proposed dynamics (19)-(20) has one zero eigenvalue related to the agreement (or consensus) on the states, i.e., all the auxiliary states reach consensus on $\mathbf{y}_{i}=0$ and all the main states reach consensus on $\mathbf{x}_{i}=\mathbf{x}^{*}$. On the other hand, the rest of eigenvalues remain in the left-half-plane which results in the stability of the proposed dynamics and convergence toward optimal point.
@@
Consider the typical example of link nonlinearity $h_{l}(\cdot)$ as log-scale quantization on the communication channels. Following Eq. (5), for sufficiently small $\rho$ one can see that $q_{l}(\cdot)$ gets arbitrarily close to the linear case. In this case, the convergence follows the approximate rate of $\mathcal{O}(\frac{1}{k})$.
@@
Uniform quantization in contrast to logarithmic quantization is not sector-bound; as shown in Fig. 1, the bounds $x\pm\frac{\rho}{2}$ do not pass through the origin. Therefore, the results on convergence and optimality do not hold for uniformly quantized data. In fact, some literature show that uniform quantization generally results in larger optimality gap, see an example for quantized resource allocation in [47]. In the next section, we clearly show by numerical simulation that log-scale quantization as compared to uniform quantization results in smaller optimality gap.
@@ IV-A Academic Example
with random parameters $a_{i,j}$ and $b_{i,j}$ in the range $[-10,10]$ such that $\sum_{i=1}^{n}\sum_{j=1}^{m}a_{i,j}=0$ and $\sum_{i=1}^{n}\sum_{j=1}^{m}b_{i,j}=0$ and $a_{i,j},b_{i,j}\neq 0$. Note that $f_{i,j}(x_{i})$ is not necessarily convex, i.e., $\nabla^{2}f_{i}(\mathbf{x})$ might be negative at some points, while the global cost summing all $f_{i,j}(x_{i})$ is strongly convex. We apply Algorithm 1 over random ad-hoc networks and structured exponential networks for both uniform and log-scale quantization. First, we compare the convergence rate and optimality gap over a time-varying exponential network (as a structured network) versus a random Erdos-Renyi (ER) network subject to log-scale quantization with $\rho=\frac{1}{128}$. The result is shown in Fig. 3. Clearly, the structured network gives a smaller optimality gap and faster convergence.
@@
Next, we compare distributed optimization over exponential graphs subject to uniform quantization versus log-scale quantization in Fig. 4 for different quantization levels $\rho$. Evidently from the figure, log-scale quantization results in a smaller optimality gap as compared to uniform quantization.
@@ IV-B Real Data-Set Example
with $\mathbf{b},c$ as the parameters of the separating hyperplane. We run and compare the distributed training over both geometric graph (as an ad-hoc network) and exponential graph (as a structured network). The optimality gap (also known as the optimization residual) in the figures is defined as $F(\overline{\mathbf{x}}^{k})-F(\mathbf{x}^{*})$ with $\overline{\mathbf{x}}^{k}=\frac{1}{n}\sum_{i=1}^{n}\mathbf{x}^{k}_{i}$.
@@
Next, we compare training over exponential networks subject to uniform quantization for different data quantization levels, see Fig. 6. As it can be observed, the uniform quantization results in a larger optimality gap in the training outcome.
Figure 6: Optimality gap of uniform quantization for distributed learning of MNIST dataset over exponential graphs.
The same comparison is performed over the geometric network and the result is shown in Fig. 7. As expected, uniform quantization results in a larger optimality gap. It is worth noting from the figures that, the optimality gap is generally larger over geometric networks as a non-structured topology.
Figure 7: Optimality gap of uniform quantization for distributed learning of MNIST dataset over geometric graphs.
@@ V-B Future Directions
Future research efforts will focus on considering the effect of quantization on different machine learning techniques, for example, distributed minimization of error-back-propagation in Neural-Networks and distributed reinforcement learning. Application to distributed optimal control is another direction of research interest.

### v3

@@ III-C Discussions
For convergence we only need $(\mathbf{1}_{n}\otimes I_{p})^{\top}H(\mathbf{1}_{n}\otimes I_{p})\succ 0$, i.e., strong-convexity of the global objective function $F(\mathbf{x})$. Thus, the local cost functions $f_{i}(\mathbf{x}_{i})$ might be non-convex, i.e., $\nabla^{2}f_{i}(\mathbf{x}_{i})$ is not necessarily positive while their summation is positive to ensure that $F(\mathbf{x})$ is strongly convex. An academic example of such a case is given in the simulation. Therefore, the convexity assumption in this paper is more relaxed as compared to many existing literature. Note that the log-scale quantization in this work is independent of Assumption 2.
@@
From Theorem 1, the proposed dynamics (19)-(20) has one zero eigenvalue related to the agreement (or consensus) on the states, i.e., all the auxiliary states reach consensus on $\mathbf{y}_{i}=0$ and all the main states reach consensus on $\mathbf{x}_{i}=\mathbf{x}^{*}$. On the other hand, the rest of the eigenvalues remain in the left-half-plane which results in the stability of the proposed dynamics and convergence toward the optimal point.
@@
Recall that one can approximate the convergence rate (or decay rate) of the linearized system dynamics given by Eq. (67) which is concluded from the dynamics (19)-(20). Following the eigen-spectrum of the system matrix $M_{q}(\alpha)$ and using Eq. (68), the convergence rate is (at least) $\exp(\max_{1\leq j\leq m}\operatorname{Re}\{{\lambda}_{2,j}\}t)$. This shows linear convergence in log-scale which is approved by the simulations in the next section. Also, for the special case of no quantization (i.e., $h_{l}(\mathbf{x}_{i})=\mathbf{x}_{i}$ and $h_{l}(\mathbf{y}_{i})=\mathbf{y}_{i}$) and for strongly convex cost functions with stochastic weight matrices $A_{\gamma},B_{\gamma}$ the convergence rate can be calculated via similar analysis as in [63, Lemmas  3.2-3.4].
@@
Uniform quantization in contrast to logarithmic quantization is not sector-bound; as shown in Fig. 1, the bounds $x\pm\frac{\rho}{2}$ do not pass through the origin. Therefore, the results on convergence and optimality do not hold for uniformly quantized data. In fact, some works in the literature show that uniform quantization generally results in a larger optimality gap, see an example for quantized resource allocation in [64]. In the next section, we clearly show by numerical simulation that log-scale quantization as compared to uniform quantization results in a smaller optimality gap.
@@ IV-A Academic Example
with random parameters $a_{i,j}$ and $b_{i,j}$ in the range $[-10,10]$ such that $\sum_{i=1}^{n}\sum_{j=1}^{m}a_{i,j}=0$ and $\sum_{i=1}^{n}\sum_{j=1}^{m}b_{i,j}=0$ and $a_{i,j},b_{i,j}\neq 0$. Note that $f_{i,j}(x_{i})$ is not necessarily convex, i.e., $\nabla^{2}f_{i}(\mathbf{x})$ might be negative at some points, while the global cost summing all $f_{i,j}(x_{i})$ is strongly convex. We apply Algorithm 1 over random ad-hoc networks and structured exponential networks for both uniform and log-scale quantization. First, we compare the convergence rate and optimality gap over a time-varying (switching) exponential network (as a structured network) versus a randomly changing Erdos-Renyi (ER) network subject to log-scale quantization with $\rho=\frac{1}{128}$. For the exponential network, we change the link weights every $100$ iterations and for the ER random network both the link weights and network structure change every $100$ iterations. The result is shown in Fig. 3. The simulation results clearly verify that the solution converges over switching networks and also, as stated in Section II-A, the structured network gives a smaller optimality gap and faster convergence.
@@
Next, we compare distributed optimization over exponential graphs under time-varying weights subject to uniform quantization versus log-scale quantization in Fig. 4 for different quantization levels $\rho$. Evidently from the figure, log-scale quantization results in a smaller optimality gap as compared to uniform quantization which verifies the statements in Section II-A. This shows the advantage of using logarithmic quantization over uniform quantization. Also, one can see that the log-scale quantization level has negligible effect on the convergence rate and optimality gap. Note that, from Remark 3, the convergence rate is linear in the log-scale. Therefore, one can approximate the convergence rate of the log-quantized case by the linear (non-quantized) solution. Also, note that considering the computational limitation of MATLAB the existing optimality gap of the log-quantized dynamics in the simulation is acceptable.
@@
Next, we compare our proposed log-quantized algorithm with some existing literature: TV-AB [65], finite-time [66], and fixed-time [67] algorithms. It should be noted that none of these algorithms are quantized and these algorithms are designed to solve strongly-convex cost functions. Therefore, we need to decrease the non-convexity parameter of the cost model (69) and set $a_{i,j}\in[-1,1]$ to make it strongly-convex for this simulation. For our log-quantized solution we set $\rho=\frac{1}{4}$. The comparison is given in Fig. 5. It is clear from the figure that our solution, although quantized, has the same performance as the non-quantized AB-type solution and shows better performance than fixed-time and finite-time solutions.
Figure 5: The comparison of the proposed log-quantized algorithm with the existing non-quantized algorithms for a strongly-convex cost model.
@@ IV-B Real Data-Set Example
with $\mathbf{b},c$ as the parameters of the separating hyperplane. We run and compare the distributed training over both geometric (as an ad-hoc network) and exponential graphs (as a structured network). The optimality gap (also known as the optimization residual) in the figures is defined as $F(\overline{\mathbf{x}}^{k})-F(\mathbf{x}^{*})$ with $\overline{\mathbf{x}}^{k}=\frac{1}{n}\sum_{i=1}^{n}\mathbf{x}^{k}_{i}$.
@@
Next, we compare training over exponential networks subject to uniform and log-scale quantization for different data quantization levels, see Fig. 7 and Fig. 8. As it can be observed, the uniform quantization results in a larger optimality gap in the training outcome.
Figure 7: Optimality gap of distributed learning of MNIST dataset over exponential graphs subject to uniform quantization with different quantization levels.
Figure 8: Optimality gap of distributed learning of MNIST dataset over exponential graphs subject to log-scale quantization with different quantization levels.
The same comparison is performed over the geometric network and the result is shown in Fig. 9 and Fig. 10. As expected, uniform quantization results in a larger optimality gap. Also, from Fig. 8 and 10 one can see that the convergence rates for different log-scale quantization rates do not change that much. It is worth noting from the figures that, the optimality gap is generally larger over geometric networks as a non-structured topology. However, the log-scale quantization adds more computational complexity to the algorithm as compared to the uniform quantization. This is because log-quantization is more precise around the origin and requires more bits for transmitting those values.
Figure 9: Optimality gap of distributed learning of MNIST dataset over geometric graphs subject to uniform quantization with different quantization levels.
Figure 10: Optimality gap of distributed learning of MNIST dataset over geometric graphs subject to log-scale quantization with different quantization levels.
@@ V-B Future Directions
One direction of future research is to more relax Assumption 1 to uniform-connectivity of the network, where the union of the graphs over a finite interval is connected. More detailed theoretical analysis and comparison between log-scale versus uniform quantization for general distributed optimization methods and quantifying the optimality gap of the discretized version for large sampling periods are set as future research. Other future research efforts will focus on considering the effect of quantization on different machine learning techniques, for example, distributed minimization of error-back-propagation in Neural-Networks and distributed reinforcement learning. Application to distributed optimal control is another direction of research interest.